. What is the speed of a typical hydrogen molecule in the star's atmosphere?
Answer #1

Here ,

Temperature , T = 7200 K

M = 0.002 Kg/mol

v_p = \sqrt { \frac{2kT}{m} } = \sqrt { \frac{2RT}{M} }

Vp = sqrt(2*8.314*7200/.002)

Vp = 7737 m/s

the most probable speed (typical speed) is 7737 m/s